### v2

# Projection-based curve pattern search for black-box optimization over smooth convex sets

Abstract
In this paper, we deal with the problem of optimizing a black-box smooth function over a full-dimensional smooth convex set. We study sets of feasible curves that allow to properly characterize stationarity of a solution and possibly carry out sound backtracking curvilinear searches. We then propose a general pattern search algorithmic framework that exploits curves of this type to carry out poll steps and for which we prove properties of asymptotic convergence to stationary points. We particularly point out that the proposed framework covers the case where search curves are arcs induced by the Euclidean projection of coordinate directions. The method is finally proved to arguably be superior, on smooth problems, than other recent projection-based algorithms and is competitive with state-of-the-art methods from the literature on constrained black-box optimization.
keywords:
Pattern search methods; Projection; Smooth constraints; Derivative-free line search
articletype: ARTICLE
{amscode}
90C56, 90C30, 90C26
1 Introduction
In this manuscript we consider the problem of minimizing a continuously differentiable function $f:C\to\mathbb{R}$ within a compact convex feasible set $C\subset\mathbb{R}^{n}$, i.e.,
$\min$ $f(x)$ (1)
s.t. $x\in C.$
More specifically, we assume that $C$ is an $n$-dimensional compact object with a smooth boundary: in other words, there exists some continuously differentiable function $g:\mathbb{R}^{n}\to\mathbb{R}$ such that $C=\{x\in\mathbb{R}^{n}\mid g(x)\leq 0\}$, its frontier is given by $\partial C=\{x\in\mathbb{R}^{n}\mid g(x)=0\}$ and there exists $x\in C$ such that $g(x)<0$, i.e., $x\in\operatorname{int}C$. While function $g$ might be unknown, we assume to have access to the Euclidean projection operator onto this set $P_{C}:\mathbb{R}^{n}\to C$ and that the cost of computing projection is sustainable.
On the other hand, while $f$ is smooth, we assume not to have direct access to its derivatives: we are thus in a setting of black-box optimization. We assume that $f$ is bounded below on $C$ by some value $f^{*}$ and that the value of $f$ cannot be computed outside the feasible set.
A vast literature has dealt with black-box optimization problems; we refer the reader to [larson2019derivative] for a recent survey on derivative-free optimization. In this paper, we put particular focus on the class of pattern search methods [hooke1961direct, lewis1999pattern, lewis2000pattern, audet2002analysis, custodio2007using, ortho09] and specifically on pattern search algorithms based on line searches. These methods, whose study has roots in a couple of pioneering works by Grippo et al. [de1984stopping, grippo1988global], includes effective algorithms to tackle constrained [Lucidi2002Derivative, lucidi2002objective] and unconstrained [lucidi2002global] black-box optimization problems, even in absence of smoothness assumptions on the objective function [fasano2014linesearch].
For the particular case of constrained problems with smooth black-box objective, which is the focus of the present work, the classical approach based on line-searches along coordinate directions (or other suitable, predefined sets of directions) has naturally be extended to deal with bound constraints [Lucidi2002Derivative]; with more complex constraints, the set of search directions shall take into account the local structure of the feasible set: while this can be reasonably done with general linear constraints, when constraints are nonlinear much more caution is required [lucidi2002objective].
A different path to handling constraints within pattern search frameworks is based on penalty approaches, either in a sequential [liuzzi2010sequential, lewis2002globally] or exact fashion [di2015derivative]: the complexity of handling constraints is moved to the objective function, which can then be optimized with methods for unconstrained problems. While effective in various settings, this latter class approaches suffer from the issue of possibly needing objective function evaluations even outside the feasible set, where the black-box might be not well defined. Moreover, this kind of methods often suffer from the high sensitivity to the choice of the penalty parameters, that heavily impact the performance of the algorithm.
Recently, a strategy was proposed in [galvan2021parameter] that overcomes these limitations under the assumption that the projection operation onto the feasible set is available: problem (1) is equivalently reformulated as an unconstrained nonsmooth problem where the objective function at any point $x$ is equal to the sum of the original objective computed at the projection of $x$ onto $C$ and the distance of $x$ from the set $C$ itself; clearly, the value corresponds to the value of the original objective when $x\in C$. The equivalent unconstrained problem can then be tackled with the CS-DFN method from [fasano2014linesearch] for nonsmooth problems.
While the projection based approach from [galvan2021parameter] has some connections to the fundamental ideas of projected gradient methods for first-order optimization [bertsekas1999nonlinear, Sec. 2.3], it cannot be actually seen as a zeroth-order adaptation of the latter. In fact, while points obtained by steps along given directions are projected onto the feasible sets in both frameworks, the projected gradient method produces a sequence of iterates all belonging to the feasible set, whereas the approach from [galvan2021parameter] possibly moves along unfeasible solutions; when this happens, it may come with some practical drawbacks: first, the points to be polled will be (almost) all unfeasible, thus requiring a high number of projections; this might result in a significant cost if projection onto $C$ is not particularly cheap. Moreover, not only the presence of a penalty part within the auxiliary objective might end up slowing down the optimization process, but its nonsmoothness also makes convergence dependent on the employment of random search directions.
The main contribution of this paper consists in a more consistent adaptation of the projected gradient method to the black-box case, under the assumption that $C$ is a full-dimensional smooth convex subset of the Euclidean space. Such an extension, which overcomes the aforementioned drawbacks of the approach from [galvan2021parameter], requires to study the behavior of searches along suitable projection arcs induced by a predefined set of search directions. To the best of our knowledge, the idea of a curve pattern search that polls the objective at points obtained, for example, by the projection of coordinate steps is novel in the literature.
It may be important to underline here that we have a different goal compared to other works, such as the recent contribution [Custódio03052024]. In the latter paper, authors suggest to emulate actual projected-gradient steps in the black-box scenario: this can be done employing a numerically approximated gradient smartly obtained using the information provided by sampled points. We rather aim to insert the projection mechanism directly in the poll step: such a technique might induce benefits like, e.g., guaranteed feasibility of sampled points. In fact, a similar strategy as the one we propose in this manuscript is mentioned (by the name DDS-proj) as a heuristic and tested in the experimental section of [Custódio03052024], either as a standalone method and as a way to improve the polling step for the simplex-gradient type approach; however, in contrast to this work, the theoretical properties of DDS-proj are not formally discussed there.
More in general, for our method there will not even be the need for a dense of search directions to achieve convergence solely via polling in constrained problems. This property does not hold true, even when $f$ is smooth, for other state-of-the-art pattern-search methods in the literature like, e.g., [ortho09] and [Custódio03052024] itself.
All in all, despite a significantly simplified structure, the algorithm we propose to use in our experiments possesses convergence guarantees and shows competitive performance w.r.t. state-of-the-art alternatives.
The rest of the paper is organized as follows: in Section 2, we review known results concerning the characterization of the feasible sets considered in this manuscript and the corresponding stationary solutions; then, we turn in Section 3 to the definition of the concept of feasible search path and the interconnection between stationarity of solutions and suitable sets of search curves. We present the novel algorithmic framework in Section 4, providing the characterization of a gradient-based version in Section 4.1 and then presenting and theoretically analyzing the derivative-free method in Section 4.2. In Section 5 we present the results of computational experiments assessing efficiency and effectiveness of the proposed method - implemented exploiting coordinate directions and projection operation - compared to another pattern search approach and a selection of state-of-the-art methods from the literature. We finally give concluding remarks in Section 6.
2 Characterization of the feasible set
As stated in the introduction, in this work we deal with feasible sets $C$ that satisfy the following assumptions:
  1. (a1)
    $C$ is convex;
  2. (a2)
    $C$ is compact;
  3. (a3)
    there exists $g:\mathbb{R}^{n}\to\mathbb{R}$ continuously differentiable such that $C=\{x\in\mathbb{R}^{n}\mid g(x)\leq 0\}$ and $g(x)<0$ for some $x\in C$.
As with any convex feasible set, we can characterize points $x$ in the feasible region by the corresponding tangent cone $T_{C}(x)$, defined as [bertsekas1999nonlinear, Def. 3.3.1]:
$T_{C}(x)=\{0\}\cup\left\{y\in\mathbb{R}^{n}\,\middle|\,\exists\{x^{k}\}\subseteq C\setminus\{x\}\text{ s.t. }x^{k}\to x\text{ and }\frac{x^{k}-x}{\|x^{k}-x\|}\to\frac{y}{\|y\|}\right\}.$
The tangent cone is always a convex cone if $C$ is convex. Moreover, $T_{C}(x)$ always contains all feasible directions at $x$, so it is easy to see that the stationarity property
$\nabla f(x)^{T}d\geq 0\quad\forall\;d\in T_{C}(x)$ (2)
is a necessary condition of optimality for problem (1). This condition in general is not immediately checkable, since we are asking for the property to hold for infinitely many directions.
However, the condition becomes numerically easy to assess if we are able to do projections onto the tangent cone and we have direct access to the gradient of the objective function.
Proposition 2.1.
A point $\bar{x}\in C$ is a stationary point in the sense of (2) for problem (1) if and only if we have
$P_{T_{C}(\bar{x})}(-\nabla f(\bar{x}))=0.$
Proof.
Assume $\bar{x}$ is stationary and let $\bar{d}=P_{T_{C}(\bar{x})}(-\nabla f(\bar{x}))$. Since $0\in T_{C}(\bar{x})$, by the properties of the projection onto the convex set $T_{C}(\bar{x})$ we have
$(-\nabla f(\bar{x})-\bar{d})^{T}(0-\bar{d})\leq 0,$
i.e.,
$\nabla f(\bar{x})^{T}\bar{d}\leq-\|\bar{d}\|^{2}.$
Since $\bar{d}\in T_{C}(\bar{x})$ by definition and $\bar{x}$ is stationary, we necessarily have $\|\bar{d}\|=0$.
On the other hand, assume $\bar{d}=P_{T_{C}(\bar{x})}(-\nabla f(\bar{x}))=0$. Let $d$ be any direction in $T_{C}(\bar{x})$. By the properties of projection we can write this time
$(-\nabla f(\bar{x})-\bar{d})^{T}(d-\bar{d})\leq 0,$
i.e.,
$0\geq(-\nabla f(\bar{x})-0)^{T}(d-0)=-\nabla f(\bar{x})^{T}d,$
which finally implies $\nabla f(\bar{x})^{T}d\geq 0$. Since $d$ is arbitrary in $T_{C}(\bar{x})$, we get the thesis. ∎
Under assumption (a3), we can go even further. Since Slater’s Constraint Qualification (CQ) is clearly verified, for every point $x\in C$ the tangent cone coincides with the cone of first order variations $V_{C}(x)$ [bertsekas1999nonlinear, Sec. 3.3.6]:
$T_{C}(x)=V_{C}(x)=\begin{cases}\{y\in\mathbb{R}^{n}\mid\nabla g(x)^{T}y\leq 0\}&\text{if }g(x)=0\text{ (i.e., }x\in\partial C\text{)},\\ \mathbb{R}^{n}&\text{otherwise},\end{cases}$
where $\nabla g(x)\neq 0$ for all $x\in\partial C$ (Slater’s CQ implies Mangasarian–Fromovitz CQ, see again [bertsekas1999nonlinear, Sec. 3.3.6]). In other words, the tangent cone is the full space for points in the interior of the feasible set, whereas it is a halfspace that varies continuously along the frontier. As noted in [lucidi2002objective, Sec. 3], when $T_{C}(x)$ is polyhedral, as in our case, we can characterize a stationary point of problem (1) by a finite number of directions. In order to do so, we first need to recall the concept of positive span of a set of directions $D=\{d_{1},\ldots,d_{m}\}$: we denote by $\operatorname{cone}(D)=\{v\in\mathbb{R}^{n}\mid\exists\,\beta\in\mathbb{R}_{+}^{m}\text{ s.t. }v=\beta_{1}d_{1}+\ldots+\beta_{m}d_{m}\}$ the cone positively spanned by the directions in $D$. We are now able to state the following proposition [lucidi2002objective, Prop. 2].
Proposition 2.2.
Let $C$ be a convex set, $\bar{x}\in C$ and assume $T_{C}(\bar{x})$ is polyhedral. Then, $\bar{x}$ is a stationary point in the sense of (2) for problem (1) if and only if
$\nabla f(\bar{x})^{T}d\geq 0\quad\forall\;d\in D,$
where $\operatorname{cone}(D)=T_{C}(\bar{x})$.
The above proposition will be crucial to tackle problems of the form (1) without having access to $\nabla f$.
3 Searching along curves
Since we will have to deal with problem (1) without the opportunity of following the gradient direction, we are interested in devising predefined patterns to explore for reducing the objective value from a given point. In the unconstrained case, any set $D$ of directions such that $\operatorname{cone}(D)=\mathbb{R}^{n}$ can be used as a basis for derivative-free line searches [lucidi2002global]; the same goes for the fortunate case of bound constraints [Lucidi2002Derivative]. In presence of more complex constraints the situation gets harder, and the set of search directions to follow has to be carefully identified for each encountered solution [lucidi2002objective]. The goal of this work is to identify a predefined search scheme that does not require to study and build a set of search directions dependent on the particular current point before polling new solutions.
To this aim, we will consider curvilinear search paths, formalized according to the following definition.
Definition 3.1.
We say that $\gamma:\mathbb{R}_{+}\to C$ is a feasible search path at $x\in C$ if $\gamma$ is a continuous curve such that $\gamma(0)=x$, $\gamma(t)\in C$ for all $t\in\mathbb{R}_{+}$ and $\gamma$ is differentiable in $t=0$.
By definition, every feasible search path $\gamma$ is associated with an initial velocity $v=\gamma^{\prime}(0)$. In the following, we will be interested in sets $\Gamma(x)$ of feasible descent paths, possibly dependent on the particular point $x\in C$, such that:
  • if $x$ is not stationary, a decrease of the objective function is attained by $\gamma(t)$ for at least one $\gamma\in\Gamma(x)$ and for values of $t$ sufficiently small;
  • if there exists $\epsilon>0$ such that $f(\gamma(t))\geq f(\gamma(0))=f(x)$ for all $t\in(0,\epsilon)$ and $\gamma\in\Gamma(x)$, then $x$ is necessarily a stationary point.
The following two propositions ensure us that a set of curves $\Gamma(x)$ does the job if the velocities of the curves positively span the tangent cone at a point, i.e., if $\operatorname{cone}(D)=T_{C}(x)$ where $D=\{v=\gamma^{\prime}(0)\mid\gamma\in\Gamma(x)\}$. Here and in the following, we will denote by $\mathcal{D}_{f\circ\gamma}$ the derivative with respect to $t$ of the composite function $f(\gamma(t))$.
Proposition 3.2.
Let $C$ be a convex set and let $\bar{x}\in C$ be a nonstationary point in the sense of (2) (i.e., $\exists\,d\in T_{C}(\bar{x}):\nabla f(\bar{x})^{T}d<0$) such that $T_{C}(\bar{x})$ is polyhedral. Let $\Gamma(\bar{x})$ be a set of feasible search paths such that $\operatorname{cone}(D)=T_{C}(\bar{x})$ with $D=\{v=\gamma^{\prime}(0)\mid\gamma\in\Gamma(\bar{x})\}$. Then, for any $\sigma>0$, there exists $\gamma\in\Gamma(\bar{x})$ and $\bar{t}>0$ such that, for all $t\in(0,\bar{t}]$, we have $f(\gamma(t))-f(\gamma(0))<-\sigma t^{2}<0$.
Proof.
By Proposition 2.2, there exists ${v}\in D$ such that $\nabla f(\bar{x})^{T}{v}<0$. Let $\gamma\in\Gamma(\bar{x})$ such that $v=\gamma^{\prime}(0)$.
Now, let us assume by contradiction that
$f(\gamma(t_{k}))\geq f(\bar{x})-\sigma t_{k}^{2}$
for a sequence $\{t_{k}\}\subseteq\mathbb{R}_{+}$ such that $t_{k}\to 0^{+}$. Rearranging and dividing both sides by $t_{k}$, we get
$\frac{f(\gamma(t_{k}))-f(\gamma(0))}{t_{k}}\geq-\sigma t_{k}.$
Taking the limits for $k\to\infty$ and recalling the definition of the sequence $\{t_{k}\}$, we get
$\frac{f(\gamma(t_{k}))-f(\gamma(0))}{t_{k}}\to\mathcal{D}_{f\circ\gamma}(0),\qquad-\sigma t_{k}\to 0,$
thus,
$\mathcal{D}_{f\circ\gamma}(0)\geq 0.$
We can then note that, by the chain rule,
$\mathcal{D}_{f\circ\gamma}(0)=\nabla f(\gamma(0))^{T}\gamma^{\prime}(0)=\nabla f(\bar{x})^{T}v.$
We therefore get that,
$\nabla f(\bar{x})^{T}v\geq 0,$
which finally gives us a contradiction. ∎
Proposition 3.3.
Let $C$ be a convex set and let $\bar{x}\in C$ such that $T_{C}(\bar{x})$ is polyhedral. Let $\Gamma(\bar{x})$ be a set of feasible search paths such that $\text{cone}(D)=T_{C}(\bar{x})$ with $D=\{v=\gamma^{\prime}(0)\mid\gamma\in\Gamma(\bar{x})\}$. Further assume that there exists $\epsilon>0$ such that $f(\gamma(t))\geq f(\gamma(0))=f(\bar{x})$ for all $t\in(0,\epsilon)$ and $\gamma\in\Gamma(\bar{x})$. Then, $\bar{x}$ is a stationary point for problem (1) in the sense of (2).
Proof.
By the assumptions, we know that for all $\gamma\in\Gamma(\bar{x})$ and for $t>0$ sufficiently small we have $f(\gamma(t))-f(\gamma(0))\geq 0$ and thus
$\frac{f(\gamma(t))-f(\gamma(0))}{t}\geq 0.$
Taking the limits for $t\to 0^{+}$, we get
$\mathcal{D}_{f\circ\gamma}(0)=\nabla f(\gamma(0))^{T}\gamma^{\prime}(0)=\nabla f(\bar{x})^{T}v\geq 0$ (3)
for all $v\in D$. Since by assumption $\operatorname{cone}(D)=T_{C}(\bar{x})$, we get the thesis from Proposition 2.2. ∎
Proposition 3.2 actually provides us not only with a guaranteed decrease result at a nonstationary point, but also ensures us that we can attain a sufficient decrease.
However, the results in Propositions 3.2 and 3.3 are clearly based on choices of $\Gamma(x)$ heavily dependent on the particular point $x$.
In the following, we are going to present a scheme, only based on the projection operator, implicitly defining curves that satisfy these suitable conditions at all points of a set $C$ satisfying assumptions (a1)-(a3). Specifically, we are going to consider, at any point $x\in C$, the paths defined by curves of the form
$\gamma_{y}(t)=P_{C}(x+ty),\qquad y\in\mathcal{Y},$ (4)
where $\mathcal{Y}$ is a suitable set of directions (not necessarily feasible nor in the tangent cone at $x$). We now start providing a preliminary result that ensures that curves of the form (4) actually define feasible search paths at $x$ with identifiable initial velocities.
Lemma 3.4.
Let $C$ be a convex set and $\bar{x}\in C$. Let $y\in\mathbb{R}^{n}$ be any direction. The curve $\gamma_{y}(t)=P_{C}(\bar{x}+ty)$ is a feasible search paths at $\bar{x}$, with an initial velocity $\gamma_{y}^{\prime}(0)=P_{T_{C}(\bar{x})}(y)$.
Proof.
The continuity of $\gamma_{y}$ straightforwardly follows from the continuity of the projection onto a convex set. Moreover, by the definition of projection, we trivially have $\gamma_{y}(t)\in C$ for all $t$ and $\gamma_{y}(0)=\bar{x}$. As for the differentiability in $t=0$, we know (see, e.g., [Shapiro2016, eq. (2)]) that $P_{C}$ is directionally differentiable at every feasible point $x\in C$, with $\mathcal{D}_{P_{C}}(x;d)=P_{T_{C}(x)}(d)$ for all $d\in\mathbb{R}^{n}$; we therefore have
$P_{T_{C}(\bar{x})}(y)=\mathcal{D}_{P_{C}}(\bar{x};y)$ $=\lim_{t\to 0^{+}}\frac{P_{C}(\bar{x}+ty)-P_{C}(\bar{x})}{t}$
$=\lim_{t\to 0^{+}}\frac{\gamma_{y}(t)-\gamma_{y}(0)}{t}=\gamma_{y}^{\prime}(0).$
At this point, what we would like to identify is a (constant) set of directions $\mathcal{Y}\subseteq\mathbb{R}^{n}$ such that we are guaranteed that, at any point $x\in C$, we have
$\operatorname{cone}(\{P_{T_{C}(x)}(y)\mid y\in\mathcal{Y}\})=T_{C}(x).$
As we are going to prove shortly, the set of coordinate directions $B=\{e_{1},\ldots,e_{n},\allowbreak-e_{1},\ldots,-e_{n}\}=\{b_{1},\ldots,b_{2n}\}$, with $e_{i}$ being the $i$-th element of the canonical basis, enjoys this property when $C$ satisfies assumptions (a1)-(a3).
Proposition 3.5.
Let $C$ be a set satisfying assumptions (a1)-(a3) and let $x\in C$. Let $D=\{v_{i}=P_{T_{C}(x)}(b_{i})\mid b_{i}\in B\}$. Then $\operatorname{cone}(D)=T_{C}(x)$.
Proof.
There are two possible cases: $x\in\operatorname{int}C$ ($g(x)<0$) or $x\in\partial C$ ($g(x)=0$). We deal with the two cases separately.
  • Let us assume $x\in\operatorname{int}C$. In this case $T_{C}(x)=\mathbb{R}^{n}$, so that $P_{T_{C}(x)}(b_{i})=b_{i}$ for any $b_{i}\in B$; then we have $D=B$ and, therefore, $\operatorname{cone}(D)=\operatorname{cone}(B)$; since $B$ is trivially a positive spanning set of $\mathbb{R}^{n}$ and $T_{C}(x)=\mathbb{R}^{n}$, the result is proven for this case.
  • Let us assume now that $x\in\partial C$. Then we have $T_{C}(x)=\{d\in\mathbb{R}^{n}\mid\nabla g(x)^{T}d\leq 0\}$ with $\nabla g(x)\neq 0$ (see Figure 1 for some visual intuition). We first show that if $\bar{d}\notin T_{C}(x)$ then $\bar{d}\notin\operatorname{cone}(D)$, i.e., $\operatorname{cone}(D)\subseteq T_{C}(x)$. Since $\bar{d}\notin T_{C}(x)$, we have $\nabla g(x)^{T}\bar{d}>0$. On the other hand, we know that $D=\{P_{T_{C}(x)}(b)\mid b\in B\}$, hence by the definition of $P_{T_{C}(x)}$ we have $\nabla g(x)^{T}v\leq 0$ for all $v\in D$. Let us assume by contradiction that $\beta_{1},\ldots,\beta_{2n}\geq 0$ exist such that $\bar{d}=\sum_{v_{i}\in D}\beta_{i}v_{i}$. We then have
    $\nabla g(x)^{T}\bar{d}=\sum_{v_{i}\in D}\beta_{i}\nabla g(x)^{T}v_{i}\leq 0,$
    which is absurd. Hence $\operatorname{cone}(D)\subseteq T_{C}(x)$.
    We now show that $T_{C}(x)\subseteq\operatorname{cone}(D)$. Let $d$ be any direction in $T_{C}(x)$. We know that $\nabla g(x)^{T}d\leq 0$. Since $\nabla g(x)\neq 0$ and $\operatorname{cone}(B)=\mathbb{R}^{n}$, we are guaranteed that there exists $b_{i}\in B$ such that $\nabla g(x)^{T}b_{i}<0$ (and thus $b_{i}\in T_{C}(x)$ and $v_{i}=P_{T_{C}(x)}(b_{i})=b_{i}$). We then have, for all $t$,
    $\nabla g(x)^{T}d$ $=\nabla g(x)^{T}(d-tb_{i}+tb_{i})$
    $=\nabla g(x)^{T}(tb_{i})+\nabla g(x)^{T}(d-tb_{i}).$
    If we set $t^{*}=\frac{\nabla g(x)^{T}d}{\nabla g(x)^{T}b_{i}}\geq 0$, we can write
    $d=p+t^{*}b_{i}=p+t^{*}v_{i},$
    with $p=d-t^{*}b_{i}$ $v_{i}\in D$ and $t^{*}\geq 0$. Moreover, $p$ is such that $\nabla g(x)^{T}p=\nabla g(x)^{T}d-\frac{\nabla g(x)^{T}d}{\nabla g(x)^{T}b_{i}}\nabla g(x)^{T}b_{i}=0$, hence $p\in H_{C}(x)=\{y\in\mathbb{R}^{n}\mid\nabla g(x)^{T}y=0\}=\partial T_{C}(x)$. Then, $d\in\operatorname{cone}(D)$ if $p\in\operatorname{cone}(D)$. Now, let $\tilde{p}(t)=p+t\nabla g(x)$; since $\tilde{p}(t)^{T}\nabla g(x)=\nabla g(x)^{T}p+t\|\nabla g(x)\|^{2}>0$ for all $t>0$, we have that $P_{T_{C}(x)}(\tilde{p}(t))=P_{H_{C}(x)}(\tilde{p}(t))$. We shall now write
    $\nabla g(x)=\sum_{i:\nabla_{i}g(x)>0}\nabla_{i}g(x)e_{i}+\sum_{i:\nabla_{i}g(x)<0}|\nabla_{i}g(x)|(-e_{i}).$
    Moreover, there certainly exist $\lambda_{1},\ldots,\lambda_{n}$ such that
    $p$ $=\sum_{i=1}^{n}\lambda_{i}e_{i}$
    $=\sum_{i:\nabla_{i}g(x)>0}\lambda_{i}e_{i}+\sum_{i:\nabla_{i}g(x)<0}(-\lambda_{i})(-e_{i})+\sum_{\begin{subarray}{c}i:\nabla_{i}g(x)=0\\ \lambda_{i}\geq 0\end{subarray}}\lambda_{i}e_{i}+\sum_{\begin{subarray}{c}i:\nabla_{i}g(x)=0\\ \lambda_{i}<0\end{subarray}}|\lambda_{i}|(-e_{i}).$
    We therefore have
    $\tilde{p}(t)=\sum_{i:\nabla_{i}g(x)>0}(\lambda_{i}+t\nabla_{i}g(x))e_{i}+\sum_{i:\nabla_{i}g(x)<0}(-\lambda_{i}+t|\nabla_{i}g(x)|)(-e_{i})$
    $+\sum_{\begin{subarray}{c}i:\nabla_{i}g(x)=0\\ \lambda_{i}\geq 0\end{subarray}}\lambda_{i}e_{i}+\sum_{\begin{subarray}{c}i:\nabla_{i}g(x)=0\\ \lambda_{i}<0\end{subarray}}|\lambda_{i}|(-e_{i}).$
    We can now observe that, for $t$ sufficiently large, $\tilde{p}(t)$ is a positive linear combination of vectors in $B$. More specifically, it is a positive linear combination of vectors in $B$ such that $\nabla g(x)^{T}b\geq 0$. In other words, letting $B_{\text{out}}=\{b\in B\mid\nabla g(x)^{T}b\geq 0\}$, for $t$ sufficiently large there exist $\beta_{1},\ldots,\beta_{|B_{\text{out}}|}\geq 0$ such that $\tilde{p}(t)=\sum_{b_{i}\in B_{\text{out}}}\beta_{i}b_{i}$. We shall now note that the projection onto the linear subspace $P_{H_{C}(x)}$ is a linear operation; we therefore get
    $P_{T_{C}(x)}(\tilde{p}(t))$ $=P_{H_{C}(x)}(\tilde{p}(t))=P_{H_{C}(x)}(p+t\nabla g(x))$
    $=P_{H_{C}(x)}(p)+P_{H_{C}(x)}(t\nabla g(x))$
    $=P_{H_{C}(x)}(p)=p,$
    and, at the same time,
    $P_{T_{C}(x)}(\tilde{p}(t))$ $=P_{H_{C}(x)}(\tilde{p}(t))=P_{H_{C}(x)}\left(\sum_{b_{i}\in B_{\text{out}}}\beta_{i}b_{i}\right)$
    $=\sum_{b_{i}\in B_{\text{out}}}\beta_{i}P_{H_{C}(x)}(b_{i})$
    $=\sum_{b_{i}\in B_{\text{out}}}\beta_{i}P_{T_{C}(x)}(b_{i}),$
    which concludes the proof since vectors $P_{T_{C}(x)}(b_{i})$ are in $D$.
Figure 1: Visualization of the construction of case $x\in\partial C$ of the proof of Proposition 3.5.
4 The algorithmic scheme
In this section we finally discuss an algorithmic approach exploiting the analysis made in Section 3. Here, we want to tackle problem (1) without accessing first-order information on the objective function. However, it is useful underlining first a derivative-based counterpart of the derivative-free method.
4.1 First-order method
An algorithmic scheme could be devised with the following update rules:
$\gamma_{k}\text{ feasible search path s.t.\ }\gamma_{k}^{\prime}(0)=P_{T_{C}(x^{k})}(-\nabla f(x^{k}))=d_{k},$ (5)
$\alpha_{k}=\max_{j=0,1,\ldots}\{\delta^{j}\Delta_{0}\mid f(\gamma_{k}(\delta^{j}\Delta_{0}))\leq f(\gamma_{k}(0))-\sigma(\delta^{j}\Delta_{0})^{2}\},$ (6)
$x^{k+1}=\gamma_{k}(\alpha_{k}),$ (7)
where $\delta\in(0,1)$ and $\Delta_{0}>0$. Algorithm (5)-(7) enjoys global convergence guarantees under a further, reasonable continuity assumption on the feasible search paths $\gamma_{k}$, as stated in the hereafter.
Assumption 1.
Let $\{x^{k}\}\subseteq C$ and let $\{\gamma_{k}\}$ be a sequence such that $\gamma_{k}$ is a feasible search path at $x^{k}$ such that $\gamma_{k}^{\prime}(0)=P_{T_{C}(x^{k})}(v_{k})$ for some sequence of directions $\{v_{k}\}$. If there exists a subsequence $K\subseteq\{0,1,\ldots\}$ such that $\lim_{k\in K,k\to\infty}x^{k}=\bar{x}$ and $\lim_{k\in K,k\to\infty}v^{k}=\bar{v}$, then there exists $\bar{\gamma}:\mathbb{R}_{+}\to C$ such that $\bar{\gamma}(t)=\lim_{k\in K,k\to\infty}\gamma_{k}(t)$ for all $t\in\mathbb{R}_{+}$ and $\bar{\gamma}^{\prime}(0)=P_{T_{C}(\bar{x})}(\bar{v})$.
Proposition 4.1.
Assume $C$ satisfies assumptions (a1)-(a3). Let $\{x^{k}\}$ be the sequence generated according to (5)-(7), assuming that $x^{0}\in C$ and that the sequence of feasible search paths $\{\gamma_{k}\}$ satisfies Assumption 1. Then the sequence $\{x^{k}\}$ admits accumulation points, each one being stationary for the problem.
The proof of this result can be found, for the sake of completeness, in Appendix A.
As already outlined in Section 3, curves $\gamma_{k}$ defined as $\gamma_{k}(t)=P_{C}(x^{k}-t\nabla f(x^{k}))$ are always feasible search paths at iterates $x^{k}$. We shall note here that, in addition, this choice of the curves satisfies Assumption 1. Indeed, if $x^{k}\to\bar{x}$, we have by the continuity of $P_{C}$ and $\nabla f$ that
$\lim_{k\to\infty}P_{C}(x^{k}-t\nabla f(x^{k}))=P_{C}(\bar{x}-t\nabla f(\bar{x}))=\bar{\gamma}(t),$
and $\bar{\gamma}^{\prime}(0)=P_{T_{C}(\bar{x})}(-\nabla f(\bar{x}))$. We can then recover the convergence properties of the projected gradient method with the line search conducted along the projection arc.
4.2 Derivative-free method
In this section we finally describe the main contribution of this paper: a derivative-free algorithm for solving problem (1) exploiting a fixed set of reference polling directions, i.e., the set $B$ of coordinate directions. The algorithm is formally described in Algorithm 1.
[…]
1: Set $k=0$ and $\tilde{\alpha}_{0}=1$. Choose $x^{0}\in C$.
2:while stopping condition not satisfied do
3:  set success = False
4:  for all $b_{i}\in B$ do
5:   Choose $\gamma_{ik}$ feasible search path at $x^{k}$ s.t. $\gamma_{ik}^{\prime}(0)=P_{T_{C}(x^{k})}(b_{i})$
6:   if $f(\gamma_{ik}(\tilde{\alpha}_{k}))\leq f(x^{k})-\sigma\tilde{\alpha}_{k}^{2}$ then
7:    $\alpha_{k}=\tilde{\alpha}_{k}$
8:    set $x^{k+1}=\gamma_{ik}({\alpha}_{k})$
9:    $\tilde{\alpha}_{k+1}=\max\{\bar{\alpha},\tau\alpha_{k}\}$
10:    set success = True
11:    break
12:   end if
13:  end for
14:  if success = False then
15:   set $\alpha_{k}=0$ and $x^{k+1}=x^{k}$
16:   set $\tilde{\alpha}_{k+1}=\delta\tilde{\alpha}_{k}$
17:  end if
18:  set $k=k+1$
19:end while
20:return $x^{k}$
Algorithm 1 Feasible search paths method
Briefly, the method at each iteration scans through a set of feasible search paths, polling points for a given tentative stepsize $\tilde{\alpha}_{k}$. As soon as one of these polling points is found providing a sufficient decrease as considered in Proposition 3.2, it is chosen as a new iterate. The tentative stepsize for the next iteration is then increased, making sure it at least jumps back to a predefined threshold when it is small. In case none of the considered feasible search paths provides sufficient decrease for the tentative stepsize, the iteration is declared unsuccessful and the tentative stepsize is reduced for the next iteration.
We begin the formal analysis of Algorithm 1 studying the asymptotic behavior of the sequences of tentative and actual stepsizes.
Proposition 4.2.
Let $K\subseteq\{0,1,\ldots\}$ be the sequence of iterations of success in Algorithm 1 and let $\bar{K}$ its complementary, i.e., the sequence of unsuccessful iterations. Let $\{\alpha_{k}\}$ and $\{\tilde{\alpha}_{k}\}$ be the sequences of stepsizes produced by Algorithm 1. The following properties hold:
  1. (i)
    $\lim\limits_{k\to\infty}\alpha_{k}=0;$
  2. (ii)
    if $K$ is infinite, then
    $\lim_{k\in K,k\to\infty}\tilde{\alpha}_{k}=0;$
  3. (iii)
    $\bar{K}$ is infinite and there exists a subsequence $K_{1}\subseteq\bar{K}$ such that
    $\lim_{k\in K_{1},k\to\infty}\tilde{\alpha}_{k}=0.$
Proof.
By the instructions of Algorithm 1, $\{f(x^{k})\}$ is nonincreasing. Hence, $\{f(x^{k})\}$ admits limit $f^{*}$ (which is finite by assumption). For all iterations $k$, we have
$f(x^{k+1})\leq f(x^{k})-\sigma\alpha_{k}^{2}.$
Note that in unsuccessful iterations it holds trivially as $\alpha_{k}=0$ and $x^{k+1}=x^{k}$. Therefore, taking the limits for $k\to\infty$, we have
$f^{*}\leq f^{*}-\sigma\lim_{k\to\infty}\alpha_{k}^{2},$
which implies that $\alpha_{k}\to 0$.
Now, let us consider the sequence $\{\tilde{\alpha}_{k}\}$. We start analyzing the subsequence $K$ of successful iterations. We can immediately note that, since $\alpha_{k}=\tilde{\alpha}_{k}$ for all $k\in K$ and $\alpha_{k}\to 0$, if $K$ is infinite, then we have
$\lim_{k\to\infty,k\in K}\tilde{\alpha}_{k}=0.$ (8)
Now, let us assume that $\bar{K}$ is finite. This would imply that $k\in K$ for all $k$ sufficiently large and thus $\alpha_{k}=$ $\tilde{\alpha}_{k}\geq\bar{\alpha}$ for all $k$. This would contradict the fact that $\alpha_{k}\to 0$.
Hence, $\bar{K}$ is infinite. If $K$ is finite, let $\bar{k}$ be the largest index in $K$. Then for all $k>\bar{k}$ we have $\tilde{\alpha}_{k}=\tilde{\alpha}_{\bar{k}+1}\delta^{k-\bar{k}-1}$, which clearly goes to zero for $k\to\infty$. In this case we would thus have $\lim_{k\in\bar{K},k\to\infty}\tilde{\alpha}_{k}=0.$
Let us now assume that both $K$ and $\bar{K}$ are infinite. Since (8) holds, for $k\in K$ sufficiently large we necessarily have $\tilde{\alpha}_{k}<\bar{\alpha}$ and thus $(k-1)\in\bar{K}$. The sequence $K_{1}\subseteq\bar{K}$ such that $(k+1)\in K$ for all $k\in K_{1}$ is thus infinite. We also have $\tilde{\alpha}_{k}=\delta^{-1}\tilde{\alpha}_{k+1}=\delta^{-1}\alpha_{k+1}$. Since $\alpha_{k}\to 0$, we then get that $\tilde{\alpha}_{k}\to 0$ for $\in K_{1}, kk\to\infty$. ∎
We also need to state another result concerning stepsizes and the sufficient decrease condition before turning to the convergence analysis.
Lemma 4.3.
There exists $\hat{\alpha}\geq\bar{\alpha}$ such that, for any unsuccessful iteration $k$ in Algorithm 1 with $k$ sufficiently large, the condition
$f(\gamma_{ik}(\hat{\alpha}\delta^{j}))-f(\gamma_{ik}(0))>-\sigma\hat{\alpha}^{2}\delta^{2j}$
holds for all $i=1,\ldots,2n$ and for all $j$ such that $\tilde{\alpha}_{k}\leq\hat{\alpha}\delta^{j}\leq\hat{\alpha}$.
Proof.
Let $m_{k}$ be the largest index such that $m_{k}<k$ and iteration $m_{k}$ is of success. We then know, by the instructions of Algorithm 1, that $\tilde{\alpha}_{k}=\tilde{\alpha}_{m_{k}+1}\delta^{k-m_{k}-1}$ and that for all $j=1,\ldots,k-m_{k}$ iteration $m_{k}+j$ was of unsuccess. Thus, for all $j=1,\ldots,k-m_{k}$, we have $x^{m_{k}+j}=x^{k}$ (which implies $\gamma_{i(m_{k}+j)}=\gamma_{ik}$), $\tilde{\alpha}_{m_{k}+j}=\delta^{j-1}\tilde{\alpha}_{m_{k}+1}$ and, for all $i=1,\ldots,2n$
$f(\gamma_{i(m_{k}+j)}(\tilde{\alpha}_{m_{k}+1}\delta^{j-1}))-f(\gamma_{i(m_{k}+j)}(0))>-\sigma\tilde{\alpha}_{m_{k}+1}^{2}\delta^{2(j-1)},$
i.e.,
$f(\gamma_{ik}(\tilde{\alpha}_{m_{k}+1}\delta^{j-1}))-f(\gamma_{ik}(0))>-\sigma\tilde{\alpha}_{m_{k}+1}^{2}\delta^{2(j-1)}.$
Since the inequality above holds for all $i=1,\ldots,2n$ and for all $j=1,\ldots,k-m_{k}$, i.e., for all $\tilde{\alpha}_{m_{k}+1}\delta^{h}$ such that $\tilde{\alpha}_{k}=\tilde{\alpha}_{m_{k}+1}\delta^{k-m_{k}-1}\leq\tilde{\alpha}_{m_{k}+1}\delta^{h}\leq\tilde{\alpha}_{m_{k}+1}\delta^{0}=\tilde{\alpha}_{m_{k}+1}$, the proof is complete if we can show that $\tilde{\alpha}_{m_{k}+1}$ is a constant greater or equal than $\bar{\alpha}$ for $k$ sufficiently large.
Now, we first consider the case where the sequence $K$ of successful iterations is infinite. By Proposition 4.2 we know $\tilde{\alpha}_{k}\to 0$ for $k\in K$, which implies that for $k\in K$ sufficiently large $\tilde{\alpha}_{k}<\frac{1}{\tau}\bar{\alpha}$ and $\tilde{\alpha}_{k+1}=\bar{\alpha}$. Therefore, since $m_{k}\to\infty$ and $m_{k}\in K$, we can deduce that for $k$ sufficiently large $\tilde{\alpha}_{m_{k}+1}=\bar{\alpha}$.
On the other hand, let $K$ be finite and let $\bar{k}$ the last successful iteration. For $k$ sufficiently large we will then have $m_{k}=\bar{k}$ and thus $\tilde{\alpha}_{m_{k}+1}=\tilde{\alpha}_{\bar{k}+1}\geq\bar{\alpha}$ by the instructions of the algorithm. ∎
We are now able to state the main convergence result of this work.
Proposition 4.4.
Assume $C$ satisfies assumptions (a1)-(a3). Let $\{x^{k}\}$ be the sequence generated by Algorithm 1, assuming the sequences of feasible search paths $\{\gamma_{ik}\}$ satisfies Assumption 1. Then the sequence $\{x^{k}\}$ produced by Algorithm 1 admits accumulation points and there exists at least one accumulation point which is stationary for the problem.
Proof.
Let us consider the sequence of unsuccessful iterations $\bar{K}$, which is infinite according to Proposition 4.2. Let us consider the further subsequence $K_{1}\subseteq\bar{K}$ such that $\tilde{\alpha}_{k}\to 0$ for $\in K_{1}, kk\to\infty$, which exists by Proposition 4.2.
By the instructions of the algorithm, $\{x^{k}\}\subseteq C$. Since $C$ is compact, $\{x^{k}\}$ admits accumulation points. We can apply the same reasoning to $\{x^{k}\}_{K_{1}}\subseteq C$ to state that there is at least an accumulation point of $\{x^{k}\}_{K_{1}}$.
Now, let us assume that any of the accumulation points of $\{x^{k}\}_{K_{1}}$ are nonstationary; in particular, let $K_{2}\subseteq K_{1}$ such that
$\lim_{k\in K_{2},k\to\infty}x^{k}=\bar{x},$
with $\bar{x}$ nonstationary, i.e., letting $\bar{d}=P_{T_{C}(\bar{x})}(-\nabla f(\bar{x}))$, $\|\bar{d}\|\geq\epsilon>0$ and thus $\nabla f(\bar{x})^{T}\bar{d}\leq-\epsilon^{2}$.
Since the iterations in $K_{2}$ are unsuccessful, we have for all $k\in K_{2}$ and for all $\gamma_{ik}$, $i=1,\ldots,2n$, that
$f(\gamma_{ik}(\tilde{\alpha}_{k}))-f(\gamma_{ik}(0))>-\sigma\tilde{\alpha}_{k}^{2}.$
Let now $q$ be an arbitrary positive integer and let $\hat{\alpha}$ be the value satisfying Lemma 4.3 . Since $\tilde{\alpha}_{k}\to 0$ for $\in K_{2}, kk\to\infty$, we have $\tilde{\alpha}_{k}<{\hat{\alpha}}\delta^{q}$ for $k\in K_{2}$ sufficiently large and then, by Lemma 4.3,
$f(\gamma_{ik}({\hat{\alpha}}\delta^{q}))-f(\gamma_{ik}(0))>-\sigma{\hat{\alpha}^{2}}\delta^{2q}.$
Dividing both sides of the inequality by ${\hat{\alpha}}\delta^{q}$ and taking the limit for $\in K_{2}, kk\to\infty$, we get
$\lim_{k\in K_{2},k\to\infty}\frac{f(\gamma_{ik}({\hat{\alpha}}\delta^{q}))-f(\gamma_{ik}(0))}{{\hat{\alpha}}\delta^{q}}\geq-\sigma{\hat{\alpha}}\delta^{q}.$
Recalling Assumption 1 for all sequences $\{\gamma_{ik}\}$, we can then write
$\frac{f(\bar{\gamma}_{i}({\hat{\alpha}}\delta^{q}))-f(\bar{\gamma}_{i}(0))}{{\hat{\alpha}}\delta^{q}}\geq-\sigma{\hat{\alpha}}\delta^{q},$
where $\bar{\gamma}_{i}$ is a feasible search path at $\bar{x}$ such that $\bar{\gamma}_{i}^{\prime}(0)=P_{T_{C}(\bar{x})}(b_{i})$.
Since $q$ is arbitrary in $\mathbb{N}$, we can take the limits for $q\to\infty$, i.e., for ${\hat{\alpha}}\delta^{q}\to 0$, obtaining
$\mathcal{D}_{f\circ\bar{\gamma}_{i}}(0)=\nabla f(\bar{x})^{T}\bar{\gamma}_{i}^{\prime}(0)\geq 0.$
By Proposition 3.5, we have $\operatorname{cone}(\{\bar{\gamma}^{\prime}_{i}(0)\mid i=1,\ldots,2n\})=T_{C}(\bar{x})$. Since $\bar{d}\in T_{C}(\bar{x})$, there exist $\beta_{1},\ldots,\beta_{2n}\geq 0$ such that $\bar{d}=\sum_{i=1}^{2n}\beta_{i}\bar{\gamma}_{i}^{\prime}(0)$; thus we can conclude that
$\nabla f(\bar{x})^{T}\bar{d}=\sum_{i=1}^{2n}\beta_{i}\nabla f(\bar{x})^{T}\bar{\gamma}_{i}^{\prime}(0)\geq 0,$
which is absurd as we had assumed $\nabla f(\bar{x})^{T}\bar{d}\leq-\epsilon^{2}<0$. Therefore, all limit points of $\{x^{k}\}_{K_{2}}$ are stationary. ∎
Remark 1.
Once again, as already outlined in Section 3, any curve of the form $\gamma_{ik}(t)=P_{C}(x^{k}+tb_{i})$ is a feasible search path at $x^{k}$. We also have that, similarly as in the gradient based case, this particular choice of the curves satisfies Assumption 1, being the limit curve, by the continuity of $P_{C}$, equal to $\bar{\gamma}_{i}(t)=P_{C}(\bar{x}+tb_{i})$ with $\bar{\gamma}_{i}^{\prime}(0)=P_{T_{C}(\bar{x})}(b_{i})$. Thanks to this property, we are finally guaranteed that the projection based method soundly achieves the desired result: a stationarity point will be reached without the need of constructing a specific set of search directions for each iterate, based on the local form of the feasible set, nor by asymptotically spanning a dense of search directions (which is required, for instance, for [ortho09, Custódio03052024, galvan2021parameter]); instead, we just need to rely on a suitable fixed set of directions and the projection operator. An example of the resulting curves is shown in Figure 2.
Figure 2: Search curves implicitly defined by the coordinate directions and projection.
5 Numerical results
In this section, we show some numerical results assessing the effectiveness and efficiency of our approach on a set of standard benchmark problems. The code [^1] The implementation code of our approach can be found at https://github.com/pierlumanzu/FSP. for the experimentation was written in Python3. The experiments were run on a computer with the following characteristics: Ubuntu 24.04 OS, Intel(R) Core(TM) i5-10600KF 6 cores 4.10GHz, 32 GB RAM.
In the rest of the section, we denote Algorithm 1 as FSP. In this experimental setting, FSP was compared with the following methods:
  • a state-of-the-art projection-based solver, proposed in [galvan2021parameter], briefly discussed in Section 1 and denoted in the following as PPM [^2] The implementation code of PPM can be found at https://github.com/jth3galv/dfppm.;
  • the approach introduced in [Custódio03052024], which integrates the Directional Direct Search (DDS) framework with the Spectral Projected Gradient (SPG) method by replacing the true gradient with a simplex-based approximation; this algorithm is denoted as SPG-DDS; in particular, we tested the same variants considered in the numerical experiments of [Custódio03052024]: SPG-DDS-proj, where infeasible poll points are projected before evaluation, and SPG-DDS-barrier, where infeasible points are discarded;
  • DDS-proj, analogous to SPG-DDS-proj but without the SPG component, which primarily differs from our method by using a dense of search directions in the poll step;
  • ORTHOMADS [ortho09], a specific instance of the Mesh Adaptive Direct Search (MADS) family of algorithms, which employs at each iteration a deterministic set of $2n$ orthogonal directions.
Regarding the stopping criteria, all algorithms had a budget of 10000 function evaluations unless explicitly stated otherwise. For FSP, we set $\delta=0.5$, $\sigma=10^{-3}$, $\tau=1.025$ and $\bar{\alpha}=10^{-6}$} . Based on preliminary experiments, we adopted a dynamic ordering of the directions in $B$: each iteration was stopped as soon as a direction proved successful, and the next iteration began with that direction. Moreover, the directions $1,\ldots,1]^{\top}$ and $1,\ldots,-1]^{\top}$ were included in $B$, and all directions were tested during the first iteration of the method. Note that these rules do not spoil convergence theory. For PPM, we set the parameters values according to the most robust configuration proposed in [galvan2021parameter], including a dynamic weighting of the distance of the point $x$ to the feasible set $C$. Both FSP and PPM were stopped when the tentative step sizes fell below a threshold of $10^{-7}$. For SPG-DDS variants and DDS-proj we adopted the parameter settings suggested in [Custódio03052024], using the Matlab code provided by the authors; for ORTHOMADS, we followed the configuration described in [ortho09] and implemented in the NOMAD library [nomad22].
Sources $n$ Problems
[abramson2008pattern, hock1980test, schittkowski2012more, gould2015cutest] 22 HS22, HS232
33 HS29, HS65
44 HS43
$6,7,8$ AS6, AS7
[gould2015cutest] 22 AKIVA, BEALE, BOXBODLS, BRANIN,
BRKMCC, BROWNBS, CAMEL6, CLIFF,
CLUSTERLS, CUBE, DANIWOODLS, BOX2
33 BARD, YFITU, ALLINIT, BIGGS3
44 DEVGLA1, HATFLDB, HIMMELBF, LEVYMONT7,
PALMER2, PALMER5D, POWERSUMB
55 DEVGLA2B, HS45, LEVYMONT8, BIGGS5
66 HART6, LANCZOS1LS
88 GAUSS1LS
$10$ HILBERTB, TRIGON2
$22$ VANDANMSLS
$25$ HATFLDC
Table 1: Benchmark of problems used for experimentation.
As for the benchmark problems, they are listed in Table 1. The first set of instances were also used in [galvan2021parameter] for testing the performance of PPM, while the second set is composed by CUTEst [gould2015cutest] problems with size $n$ ranging from 2 to 25. Unless otherwise stated, all the instances were tested using a unit hyper-sphere constraint, i.e., $C=\{x\in\mathbb{R}^{n}\mid\|x-c\|^{2}\leq 1\}$, where $c\in\mathbb{R}^{n}$ denotes the center of the hyper-sphere. We recall that the projection onto such set can be calculated in closed form according to:
$P_{C}(x)=\begin{cases}c+\frac{x-c}{\|x-c\|},&\mbox{if }x\not\in C,\\ x,&\mbox{otherwise.}\end{cases}$
Both algorithms start, in each test problem, with the same feasible solution.
To assess the performance of the algorithms, we primarily focused on the number of function evaluations $n_{f}$ required by each method, i.e., the typical efficiency metric in black-box optimization; we are also interested in the number $n_{p}$ of times the projection of an unfeasible point is computed; analyzing this quantity provides an insight about efficiency in cases where projection is available, but costly; we also looked the objective function value $f^{\star}$ returned by each method. For a compact visualization of the results, we made use of the performance profiles [dolan2002benchmarking] and the data profiles [dataprof09], which represent cumulative distribution functions showing, for each solver, the percentage of problems solved within a tolerance $\varepsilon>0$ given a specific budget of function evaluations.
5.1 First set of problems
In this section, we discuss the results reported in Tables 2-3, obtained by FSP and PPM on the first set of problems of Table 1 with hyper-sphere centers $c=[0,\ldots,0]^{\top}$ and $c=[5,\ldots,5]^{\top}$, respectively.
Problem FSP PPM
$f^{\star}$ $n_{f}$ $n_{p}$ $f^{\star}$ $n_{f}$ $n_{p}$
HS22 1.528 241 128 1.528 327 256
HS232 -0.038 206 109 -0.038 434 321
HS29 -0.192 193 97 -0.192 365 261
HS65 26.548 440 246 26.548 553 469
HS43 -21.435 665 365 -21.435 519 444
AS6 ($n=6$) 2.101 351 177 2.101 891 722
AS6 ($n=7$) 2.708 402 203 2.708 1314 1088
AS6 ($n=8$) 3.343 451 227 3.343 3754 3424
AS7 ($n=6$) 0.0 1047 26 0.0 313 4
AS7 ($n=7$) 0.0 1336 31 0.0 364 4
AS7 ($n=8$) 0.0 1628 38 0.0 415 4
Table 2: Results on the first set of problems listed in Table 1 with hyper-sphere centered at $c=[0,\ldots,0]^{\top}$.
Problem FSP PPM
$f^{\star}$ $n_{f}$ $n_{p}$ $f^{\star}$ $n_{f}$ $n_{p}$
HS22 16.0 242 129 16.0 336 259
HS232 -29.373 234 133 -29.373 2037 2034
HS29 -173.494 202 102 -173.494 831 822
HS65 0.0 438 16 0.0 336 5
HS43 -12.436 539 303 -12.436 571 483
AS6 ($n=6$) 77.404 337 176 77.404 1822 1639
AS6 ($n=7$) 91.834 385 201 91.834 6883 6780
AS6 ($n=8$) 106.373 433 226 106.373 1727 1529
AS7 ($n=6$) 126.505 337 176 126.505 1023 823
AS7 ($n=7$) 149.542 385 201 149.542 830 623
AS7 ($n=8$) 172.716 433 226 172.716 733 488
Table 3: Results on the first set of problems listed in Table 1 with hyper-sphere centered at $c=[5,\ldots,5]^{\top}$.
We immediately observe that the two algorithms always returned solutions with the same function value. Moreover, an overall advantage of FSP w.r.t. its competitor is attested in the table in terms of both $n_{f}$ and $n_{p}$. These results are particularly encouraging if we take into account that PPM is based on CS-DFN [galvan2021parameter], which implements some sophisticated mechanisms aimed at speeding up the computation (such as direction-specific stepsizes, extrapolation steps), compared to the very basic setup of FSP. By a careful look at the results, we also get some insight from the cases of AS7 instances with $c=[0,\ldots,0]^{T}$ and HS65 instance with $c=[5,\ldots,5]^{T}$, where PPM has a better performance: in these problems the solution lies in the interior of the unit hyper-sphere. We can speculate that the refined mechanism of CS-DFN have a greater impact in this “unconstrained” scenario; yet, FSP is still able to solve the problem with a reasonable cost; a future integration of the CS-DFN mechanisms within FSP might thus be worth of future investigations.
Problem FSP PPM
$f^{\star}$ $n_{f}$ $n_{p}$ $T$ $f^{\star}$ $n_{f}$ $n_{p}$ $T$
HS29 -22.627 406 221 2.77 -22.627 634 565 4.92
Table 4: Results on the original HS29 problem.
We conclude the section comparing the two approaches on the original formulation of the HS29 problem, whose ellipsoidal constraint respects the feasible set assumptions made in Section 2. In this case, the projection onto the feasible convex set $C$ is not available in closed form: the use of a solver is required (we employed the CVXOPT [^3] https://cvxopt.org software package) and the projection operation effectively becomes a significant element of the computational cost of the algorithm. The results are reported in Table 4. We get similar numbers as for HS29 with the unit sphere constraint: both the algorithms reached the same optimal value and FSP outperformed PPM on all the other metrics. In particular, we can observe that FSP performed about 1/2 of the projections of PPM, and such result is also reflected on the overall elapsed time ($T$) spent by the two algorithms.
5.2 Cutest problems
In this second experimental section, we first compared FSP and PPM on the CUTEst problems reported in Table 1 under both types of unit hyper-sphere constraints, i.e., $c=[0,\ldots,0]^{\top}$ and $c=[5,\ldots,5]^{T}$. For a compact view of the results, we report in Figure 3 the performance profiles achieved by the two approaches in terms of $n_{f}$ and $n_{p}$. We shall remark that both the algorithms managed to achieve the same optimal value on all the tested instances; we do not report the numbers here for the sake of brevity.
Figure 3: Performance profiles in terms of $n_{f}$ and $n_{p}$ obtained by FSP and PPM on the Cutest problems listed in Table 1 under both types of unit hyper-sphere constraints, i.e., $c=[0,\ldots,0]^{T}$ and $c=[5,\ldots,5]^{T}$.
Again, FSP proved to consume less function evaluations overall w.r.t. the competitor. Such performance gap is even more accentuated if we look at the number of projections $n_{p}$.
Finally, we compared our method against the SPG-DDS variants, DDS-proj and ORTHOMADS on the benchmark set, using a budget of 25000 function evaluations. Given the structural differences between the solvers, we utilized data profiles for a concise comparison. Figure 4 shows these profiles for tolerances $\varepsilon\in\{10^{-2},10^{-4},10^{-6},10^{-8}\}$ against the SPG-DDS variants and DDS-proj, while Figure 5 shows the same profiles against ORTHOMADS.
We first observe that SPG-DDS-barrier performed significantly worse than our method, SPG-DDS-proj and DDS-proj. This provides empirical evidence that projection of infeasible points before polling is crucial for obtaining an effective procedure. For accuracy levels $\varepsilon\in\{10^{-2},10^{-4},10^{-6}\}$, our method was competitive with SPG-DDS-proj and DDS-proj. This arguably indicates that both the use of a dense of search directions and the eventual use of simplex-gradient steps yields no clear advantage over the simpler strategy used within our method. When increasing the required accuracy to $\varepsilon=10^{-8}$, our method even proved to be the most effective choice, outperforming all three competitors.
In the comparison with ORTHOMADS, both methods exhibited similar performance at the lowest accuracy level ($\varepsilon=10^{-2}$). However, as the accuracy requirement increased, our method achieved substantially better results. ORTHOMADS employs a dense of search directions as well, with an extreme-barrier strategy for handling constraints; this combination did not allow it to match the performance of our method, which relies only on coordinate directions.
Figure 4: Data profiles obtained by FSP, SPG-DDS-proj, SPG-DDS-barrier and DDS-proj for tolerances $\varepsilon\in\{10^{-2},10^{-4},10^{-6},10^{-8}\}$ on the Cutest problems listed in Table 1, under unit hyper-sphere constraints with both $c=[0,\ldots,0]^{T}$ and $c=[5,\ldots,5]^{T}$.
Figure 5: Data profiles obtained by FSP and ORTHOMADS for tolerances $\varepsilon\in\{10^{-2},10^{-4},10^{-6},10^{-8}\}$ on the Cutest problems listed in Table 1, under unit hyper-sphere constraints with both $c=[0,\ldots,0]^{T}$ and $c=[5,\ldots,5]^{T}$.
6 Conclusions
In this paper, we focused on black-box optimization problems where the objective function is a smooth function with inaccessible derivatives, whereas the feasible set is a smooth, convex closed set. We introduced the concept of feasible search path, i.e., a curve contained in the feasible set, starting at a feasible solution, with suitable regularity properties. We discussed the properties required by sets of such curves not only to be sufficient for the characterization of stationarity, but also for guaranteeing the decrease of the objective function if a backtracking search is conducted from a nonstationary point. Then, we presented a pattern search algorithm that polls points along feasible search paths, showing that it is provably convergent to stationary points, without resorting to asymptotically dense sets of directions. Of particular interest is the special case, fitting the general framework, of search paths defined by the projection of steps carried out along coordinate directions. The corresponding algorithm is shown to be computationally competitive for this class of problems with a selection of state-of-the-art methods from the literature. Worthy of special attention are the comparisons with other projection-based derivative-free approaches: one transforms the original problem into a nonsmooth unconstrained one, the other mixes (without proven convergence guarantees) our mechanism with an ORTHOMADS-type selection of polling directions and a simplex-gradient approach. Future work might focus on the extension of this contribution to more general convex sets of constraints.
Declarations
Funding
No funding was received for conducting this study.
Disclosure statement
The authors report there are no competing interests to declare.
Code availability statement
The implementation code of the approach presented in the paper can be found at https://github.com/pierlumanzu/FSP.
Notes on contributors
Xiaoxi Jia received her M.Sc. degree in mathematics (Optimization) from Nanjing Normal University in China in 2020. She then obtained her Ph.D. degree in Optimization in 2023 from the University of Wuerzburg in Germany. She has completed a one-year and one-month postdoctoral research at Saarland University in Germany. Her main research focuses on nonconvex and nonsmooth optimization problems.
Matteo Lapucci received his bachelor and master’s degree in Computer Engineering at the University of Florence in 2015 and 2018 respectively. He then received in 2022 his PhD degree in Smart Computing jointly from the Universities of Florence, Pisa and Siena. He is currently Assistant Professor at the Department of Information Engineering of the University of Florence. His main research interests include theory and algorithms for sparse, multi-objective and large scale nonlinear optimization.
Pierluigi Mansueto received his bachelor and master’s degree in Computer Engineering at the University of Florence in 2017 and 2020, respectively. He then obtained his PhD degree in Information Engineering from the University of Florence in 2024. Currently, he is a Postdoctoral Research Fellow at the Department of Information Engineering of the University of Florence. His main research interests are multi-objective optimization and global optimization.
ORCID
Xiaoxi Jia: 0000-0002-7134-2169
Matteo Lapucci: 0000-0002-2488-5486
Pierluigi Mansueto: 0000-0002-1394-0937
Acknowledgements
The authors are very grateful to Prof. M. Sciandrone and Dr. T. Trinci for the fruitful discussions, and to Prof. A. L. Custódio for sharing the SPG_DDS code. We would like to thank the two anonymous reviewers for their valuable comments, which helped us significantly improve the quality of the manuscript.
Appendix A Proof of Proposition 4.1
Proof.
By the instructions (5)-(7), $\{x^{k}\}\subseteq C$. Since $C$ is compact, accumulation points exist.
Now, let us assume that there exists a subsequence $K\subseteq\{0,1,\ldots\}$ such that
$\lim_{k\in K,k\to\infty}x^{k}=\bar{x},$
with $\bar{x}$ nonstationary, i.e., letting $\bar{d}=P_{T_{C}(\bar{x})}(-\nabla f(\bar{x}))$, $\|\bar{d}\|\geq\epsilon>0$ and thus $\nabla f(\bar{x})^{T}\bar{d}\leq-\epsilon^{2}$.
By the instructions of the algorithm we also know that the entire sequence $\{f(x^{k})\}$ is nonincreasing: it thus admits limit $f^{*}$, which is finite since $f$ is bounded below. Moreover, by the sufficient decrease condition we have
$f(x^{k+1})-f(x^{k})=f(\gamma_{k}(\alpha_{k}))-f(\gamma_{k}(0))\leq-\sigma\alpha_{k}^{2}\leq 0.$
Taking the limits for $\in K, kk\to\infty$, we immediately get that
$\lim_{k\in K,k\to\infty}\alpha_{k}=0.$
Hence, for any $q\in\mathbb{N}$, we have for $k\in K$ sufficiently large that the step $\delta^{q}\Delta_{0}$ does not satisfy the sufficient decrease condition, i.e.,
$f\left(\gamma_{k}\left(\delta^{q}\Delta_{0}\right)\right)-f(\gamma_{k}(0))>-\sigma(\delta^{q}\Delta_{0})^{2}.$
Dividing both sides of the above inequality by $\delta^{q}\Delta_{0}$ and taking the limit for $\in K, kk\to\infty$, we then get
$\lim_{k\in K,k\to\infty}\frac{f\left(\gamma_{k}\left(\delta^{q}\Delta_{0}\right)\right)-f(\gamma_{k}(0))}{\delta^{q}\Delta_{0}}\geq-\sigma\delta^{q}\Delta_{0}.$
We also know that $\gamma^{\prime}_{k}(0)=P_{T_{C}(x^{k})}(-\nabla f(x^{k}))$ for all $k$ and that $\nabla f(x^{k})\to\nabla f(\bar{x})$ as $x^{k}$ goes to $\bar{x}$. Recalling Assumption 1 we can then write
$\frac{f\left(\bar{\gamma}\left(\delta^{q}\Delta_{0}\right)\right)-f(\bar{\gamma}(0))}{\delta^{q}\Delta_{0}}\geq-\sigma\delta^{q}\Delta_{0}.$
Since $q$ is arbitrary in $\mathbb{N}$, we can take the limit for $q\to\infty$, i.e., for $\delta^{q}\Delta_{0}\to 0$, obtaining
$\mathcal{D}_{f\circ\bar{\gamma}}(0)=\nabla f(\bar{\gamma}(0))^{T}\bar{\gamma}^{\prime}(0)\geq 0.$
Recalling again Assumption 1, we have $\bar{\gamma}^{\prime}(0)=P_{T_{C}(\bar{x})}(-\nabla f(\bar{x}))=\bar{d}$ and so we finally get
$\nabla f(\bar{x})^{T}\bar{d}\geq 0,$
which is absurd as we had assumed $\nabla f(\bar{x})^{T}\bar{d}\leq-\epsilon^{2}<0$.